A 50 kg person steps on a scale while holding a 5 kg cat. What does the scale read in pounds? (Assume 1 pound equals 4.5 N
1
Expert's answer
2019-04-11T09:39:54-0400

Total weight:


"W=(M+m)g=(50+5)9.8=539 N."

The scale reads:


"W= \frac{539}{4.5}=120 pounds."